OpenStudy (anonymous):

Does anyone know of a reference list of math formulas in LaTeX format? NOT symbols. NOT syntax. NOT manual. A library or a collection of frequently used mathematical formulas, functions, identities, properties, etc. Things like "tanx=sinx/cosx", "a*0=0", "De^x=e^x".

OpenStudy (anonymous):

No, not a manual. A list of math formulas in LaTeX format. Things like \tan {x} = \frac {\sin {x}} {\cos {x}} without having to type this out myself.

Parth (parthkohli):

Dude! Equation editor xP

sam (.sam.):

That you have to type it yourself, but some of it is here http://en.wikipedia.org/wiki/Help:Displaying_a_formula

sam (.sam.):

\[\curvearrowleft \circlearrowleft \Lsh \upuparrows \rightrightarrows \rightleftarrows \rightarrowtail \looparrowright\]Really works

sam (.sam.):

\[\bigstar \cup, \Cup, \sqcup, \bigcup, \bigsqcup, \uplus, \biguplus \bigoplus, \bigotimes, \bigodot\]

OpenStudy (anonymous):

There gotta be a pre-existing library of frequent functions somewhere. @ParthKohli Do you know any good ones?

Parth (parthkohli):

http://en.wikipedia.org/wiki/Help:Displaying_a_formula#Functions.2C_symbols.2C_special_characters We have all possible symbols here in \( L^{A}T_{E}X \)

Parth (parthkohli):

What symbol do you want to display?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow \alpha \beta \gamma }\)

Parth (parthkohli):

Also remember, that there is a Latex practicing group

OpenStudy (anonymous):

I am not after symbols or syntax. I am after frequently used formulas. If I want to enter def for tan \[\tan {x} = \frac {\sin {x}} {\cos {x}} \] I don't want to type if out by hand, I want to copy-paste or macro it in.

OpenStudy (anonymous):

\# \$ \% \textasciicircum{} \& \_ \{ \} \~{} \textbackslash{}

OpenStudy (anonymous):

lazy :P @sakh

OpenStudy (anonymous):

\textstyle \int\limits_{-N}^{N} e^x\, dx

OpenStudy (anonymous):

\[\bigcup_{i=_1}^n E_i \]

OpenStudy (anonymous):

\[\left ( \frac{1}{2} \right ) \]

OpenStudy (anonymous):

\[\left / \frac{a}{b} \right \backslash \]

OpenStudy (anonymous):

\[\big( \Big( \bigg( \Bigg( \dots \Bigg] \bigg] \Big] \big]/ \]

OpenStudy (anonymous):

\[\boldsymbol{\alpha}\]

Parth (parthkohli):

@Rohangrr please don't post inappropriate answers.

OpenStudy (anonymous):

\[\mathcal{A} \mathcal{B} \mathcal{C} \mathcal{D} \mathcal{E} \mathcal{F} \mathcal{G}\] is given when you thpe: \mathcal{A} \mathcal{B} \mathcal{C} \mathcal{D} \mathcal{E} \mathcal{F} \mathcal{G}

OpenStudy (anonymous):

\mathfrak{A} \mathfrak{B} \mathfrak{C} \mathfrak{D} \mathfrak{E} \mathfrak{F} \mathfrak{G}

OpenStudy (anonymous):

\[\mathfrak{A} \mathfrak{B} \mathfrak{C} \mathfrak{D} \mathfrak{E} \mathfrak{F} \mathfrak{G}\]

OpenStudy (anonymous):

pretty much just google what you want and it will show up. there are tons of lists of symbols

OpenStudy (turingtest):

here's a very simple list that used when I first started learning latex: http://omega.albany.edu:8008/Symbols.html it's certainly not a complete list though, so if you want to know how something is done, right-click it-->show as-->tex commands and you will see the formula in latex

OpenStudy (anonymous):

Thanks all, but I am not after symbols. I am after a library of functions.

OpenStudy (anonymous):

i think it is not clear what you are after. a list of functions?

OpenStudy (turingtest):

what do you mean by a library of functions? like the quadratic formula, and Maxwell's eqns all ready to go in latex?

OpenStudy (turingtest):

...cuz if you know latex you can type pretty much any formula...

sam (.sam.):

@TuringTest yeah its like the quadratic formula and Maxwell's eqns,

OpenStudy (turingtest):

it's much easier to just learn a little latex than to find the formulas pre-typed for you but you're not the asker anyway @.Sam. :P

sam (.sam.):

I can literally read his mind :P , lol

OpenStudy (turingtest):

I see...

sam (.sam.):

He said already QUOTE: " I am not after symbols or syntax. I am after frequently used formulas. If I want to enter def for tan tanx=sinxcosx I don't want to type if out by hand, I want to copy-paste or macro it in. "

OpenStudy (kymber):

@ParthKohli There's a Latex practicing group?

Parth (parthkohli):

Yep @kymber

OpenStudy (kymber):

On OpenStudy? I don't see it

OpenStudy (kymber):

Oh :D Thanks!

Parth (parthkohli):

It's created by a user, not an official group/subject

OpenStudy (turingtest):

well I stand by suggestion that it is much easier to learn the basic latex and type it by hand; it's not that hard

sam (.sam.):

@TuringTest that's what I'm thinking too, typing the equations out its not that hard :)

OpenStudy (turingtest):

once you learn latex, write the formulas you want out yourself, then copy paste the syntax on a word document or something for future copy-paste availability.

OpenStudy (turingtest):

...then you have our own custom list of formulas

OpenStudy (turingtest):

I should probably do that lol so tired of typing the quadratic formula

Parth (parthkohli):

I just copy the basic latex footprint and paste it whenever I want.

Parth (parthkohli):

I always have a notepad file opened up

OpenStudy (anonymous):

if you need a text editor there are plenty

OpenStudy (anonymous):

i use something called texmaker, works great in linux but i hear it is good in windoze too

Parth (parthkohli):

No thanks satellite, notepad is my personal preference. It's for geeks :D

OpenStudy (anonymous):

yeah right. i used to roll my own cigarettes too, but i didn't cure the tobacco and mill my own paper

OpenStudy (anonymous):

Sam's right, its exactly what I want. It's not hard, just redundant.

OpenStudy (kinggeorge):

This is a nice reference that I refer to quite often. http://www.artofproblemsolving.com/Wiki/index.php/LaTeX:Symbols

OpenStudy (kinggeorge):

Another place that has some excellent examples. http://en.wikibooks.org/wiki/LaTeX/Mathematics